• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kai / Diancheng/ Alex. Camson

    Log entry time 09:44:13 on November30,2009

    Entry number 303677


    keyword=Blumi HV was stuck at #14335,#14336. I will still keep them as "production".

    #14335,#14336 are two production runs right after the Al dummy test runs. The raw Blumi signals in these two runs are abnormal. Some are too large, and some others even go to the negative side (Fig1 is the asy signal.). We checked the HV gui and found that the HV for BLumi monitors were stuck there without any response. We close and restart the hvgui. The HV display looks fine now (Fig2), but we will still keep an eye on how it goes in the next production run after the current access. Except for the blumi, other data looks normal than other production runs. Therefore, my opinion is that it is because of the HV malfuction that leads to the Lumi monitor display problem, and I will still keep them as production runs.

    FIGURE 1

    FIGURE 2